Calculate elasticities for the following cases with appropriate names. Identify what types
of goods they are, and comment on the value of the elasticity, i.e., elastic or inelastic:

a) A 10 percent increase in the price of cars reduces the quantity demanded for cars by 4
percent.

b) A 10 percent increase in the gasoline price reduces the demand for new cars by 4 percent.

c) A 10 percent increase in the income reduces the demand for used cars by 4 percent.
1
Expert's answer
2020-10-29T07:23:06-0400

Solution

(A) price elasticity of demand =

%change in quantity demanded/ % change in price

So

Price elasticity of demand ="\frac{4}{10}"

"=0.4"

Value of elasticity <1

So it is Inelastic demand . in this case Cars are substitute goods.

(B)

Cross elasticity of demand

"=\frac{4}{10}\\=0.4"

And cross elasticity of demand is < 1


So it's Inelastic demand.

Gasoline and car are complementry goods because if gasoline will consumed low then car be also consumed low and vice versa.

(C)

Income elasticity of demand

"=\frac{4}{10}"

=0.4

0.4<1 so Income elasticity of demand is less then one so it's inelastic demand. It's substitution effect .